OpenStudy (anonymous):

A sheet of lead reduces the intensity of a beam of gamma-rays by a factor 0.4. Find the half value thickness of the lead sheet which will reduce the intensity to half of its initial value.

OpenStudy (anonymous):

Let the intensity of the beam of gamma rays=Io hence I=Io*e^(-n*x)).....where n is refractive index of the lead. and x is the thickness of it. =>I/Io=e^(-n*x) =>0.4=e^()

OpenStudy (anonymous):

=>0.4=e^(-n*x)..(i) similarly 0.5=e^(-n*x)...(ii) from (i) and (ii) find x. and the half value of thickness=x/2 .find it...